A football was kicked at an angle of 450 with an initial velocity of 25.7m/s.

a. What was the horizontal displacement of the ball?
1
Expert's answer
2020-12-14T07:17:34-0500

The horizontal displacement is


"R=\frac{v^2\text{ sin}(2\theta)}{g}=\frac{25.7^2\text{ sin90°}}{9.8}=67.4\text{ m}."